OpenStudy (anonymous):

calculate step by step, if existent, via l'Hopital rule: lim-->2 ln(x^2-3)/x^2-4

myininaya (myininaya):

derivative of top/ derivative of bottom

OpenStudy (anonymous):

i have difficulty with finding the derivative of top ln(x^2-3)

OpenStudy (nubeer):

Hint ln(x+1) [1/(x+1) ] . d/dx (x+1)

OpenStudy (anonymous):

well, what will be the answer if i go step by step?

OpenStudy (nubeer):

i dont know.. why dont you solve and put steps here .. if you do something wrong i wil let u know

OpenStudy (richyw):

\[\frac{d}{dx}\left[\ln\left(x^2-3\right)\right]\]You just need to use the chain rule\[\frac{1}{x^2-3}\cdot\frac{d}{dx}\left(x^2-3\right)=\frac{2x}{x^2-3}\]I already showed you how to check wolfram alpha first for step-by step solutions.

OpenStudy (anonymous):

thank you so much... i was confused whether i can use chain or quotient rule in l'hopital or not

OpenStudy (richyw):

You can use whatever you want to figure out the derivatives. All that matters is the derivative exists (and of course you have an appropriate indeterminate form)